OpenStudy (anonymous):

Please HELP! 3x+6/8 = 7x-1/6

OpenStudy (anonymous):

\[\frac{ 3x+6 }{ 8 } = \frac{ 7x-1 }{ 6 }\]

OpenStudy (anonymous):

Okay, so I am not looking for the answer. I am just confused on how to do this with going through the steps. Will someone please explain to me what I need to do? Thanks so much. I guess I'm solving for x, by the way.

OpenStudy (anonymous):